OpenStudy (anonymous):

I need help desperately! please show the steps involved with an explanation I am so lost and want to understand! 1- A box is to be constructed from a sheet of cardboard that is 20 cm by 60 cm, by cutting out squares of length x by x from each corner and bending up the sides. What is the maximum volume this box could have? 2- Find the amplitude and period of y= -3cos(2x+3). use your calculator to graph the function and state its symmetry. find the first positive x-intercept using your calculator's zero function. 3-find two functions f(x) and g(x) such that f[g(x)]= x but g(f(x)]

OpenStudy (anonymous):

does not equal x. 4- state the vertical , horizontal asymptotes and zeros of the rational f(x)= x^2+3x+2/x^2+5x+4. why is there no zero at x=-1? 5-give an example and explain why a polynomial can have fewer x-intercepts than its number of roots.

OpenStudy (rational):

chances are very less for ppl to help you when you post more than one question in the same thread

OpenStudy (sepeario):

especially because many people (like myself) don't know how to solve these type of difficult questions :(

OpenStudy (anonymous):

awok ty anyway.

OpenStudy (sparrow2):

v=(60-2x)(20-2x)x then you need a derivate of this function and than derivaitive=0

OpenStudy (anonymous):

 put dv/dx=0 and find values of x find d^2v/dx^2 at each point and find where d^2v/dx^2 is negative .